OpenStudy (anonymous):

How do I solve this? x^2-10x=0

OpenStudy (saifoo.khan):

Find the GCF first, (or in other words, find the common term).

OpenStudy (saifoo.khan):

Here, the common thing among these two is "x" x(x-10) = 0

OpenStudy (saifoo.khan):

Now, set them equal to zero. x = 0 x - 10 = 0 x = 0 x = 10

OpenStudy (mimi_x3):

\[\large x(x-10) = 0 \] \[\large x = 0 , 10\]

OpenStudy (anonymous):

x = 0 , 10

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (anonymous):

U r welcome

OpenStudy (dumbcow):

haha thats a lot of medals for this problem

OpenStudy (anonymous):

That's natural!!!!!!!!

OpenStudy (saifoo.khan):

Not much only 19. ;)